How much interest will I have to pay if I borrow $6000 at an annual rate of 13% for a period of 8 months?
Answer by ikdeep(226)

Here we have to calulate the interest,,
SO,let the interest = I
Principle = $6000 (given in question)
Rate of interest = 13% (again given in question)
Time = 8 months
Now here as we can't calculate interest on monthly basis, so lets 1st convert 8 months into years .....and it goes like this .....
12 months = 1 year
1 month = 1/12 year
8 months = 1/12 *8...this gives
= 2/3 year
Now, according to the formulae,,,,,
interest = (Principle * rate of interest * time)/100
For ease,,,just learn it as ...I = PRT/100...
Now put the values of principle ,,rate of interest and time (which we calculated) from above in the formulae.....we get...
I = (6000 * 13 * 2/3)/100
= (2000*13*2)/100
= (52,000)/100
= $520

Therfore, Interest = $520